In summer, a given sample of milk turns sour at 27°C in 48 hours. In a 2 refrigerator at 2°C, milk can be stored three times longer before it sours. Calculate the time taken by the milk to sour at 37 degree celsius

Grade:9

1 Answers

Arun
6 years ago
logk2/k1=Ea/2.303*R(T2-T1/T1*T2)log1/3=Ea/2.303*8.314(275-300/275*300)Ea=30.14kjsince the rate at 27degree clsius is 3 times more faster than at 2 degree celsius.Hope it helps